What horizontal force P can be applied to a 100-kg block in a level surface with coefficient of friction of 0.2, that will cause and acceleration of 2.50 m/s2?


1
Expert's answer
2022-05-18T15:27:53-0400
"ma=F-F_f\\
ma=F-\mu mg"

"F=m(a+\mu g)\\
={100\:\rm kg*(2.50\:\ m/s^2+0.2*9.8\:m /s^2)}\\
=446\:\rm N"